OpenStudy (anonymous):

A rectangle has a base of 5 inches and a height of 7 inches. If the dimensions are cut in half, what will happen to the area of the rectangle? It will be half of the original size. It will be the same as the original size. It will be four times the original size. It will be one-fourth the original size.

ganeshie8 (ganeshie8):

make a guess

OpenStudy (anonymous):

c

OpenStudy (jjwalton04):

Derp. Well no, heres the explination It will be 1/4 the original size. original size 5"*7"=35sq " 5/2 to divide base in half = 2 1/2" 7/2 to divide height in half = 3 1/2" then multiple 2 1/2" x 3 1/2" first change mixed numbers to improper fractions 2 1/2 = 5/2 and 3 1/2=7/2 5/2*7/2=35/4 or 8 3/4 sq in. So it is 1/4 the original size you can check this by multiplying 35/4*4/1 = 35 square inches the original was 35 square inches

OpenStudy (jjwalton04):

Explanation

ganeshie8 (ganeshie8):

really ? how can you ever get more area by reducing the dimensions ?

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

cut width in half also cut length in half

OpenStudy (anonymous):

a

ganeshie8 (ganeshie8):